• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:12:38 on June12,2002

    Entry number 85939

    keyword=Shift Summary

    Down for machine development all day.
    Ran tests of software controls for right arm and found no problems.
    Tech's did some work on right dipole power supply (removed UPS that was believed to be blowing fuses).
    Tech's also working on lead flow readback for left Q1. Lead flow is fine, interlock is fine, but readback number was not (was giving negative leadflow).
    At end of shift right dipole is ramping back up to 2.8607 GeV/c.
    Scott is still working on left Q1 lead flow. (It's at least positive now but not yet calibrated).
    We've been on controlled access all day with several people going in and out of the Hall. The next shift leader, Pete, is reminded to fill out the controlled access checklist before securing from controlled access.